OpenStudy (anonymous):

how do i sketch the curve for (x^2+y^2)^3=4x^2y^2 ?

OpenStudy (anonymous):

it is with polar coordinates?

OpenStudy (dumbcow):

using polar coordinates @ = theta x=rcos@ y=rsin@ x^2 + y^2 = r^2 substituting this in (r^2)^3 = 4(r^2(cos@)^2)(r^2(sin@)^2) r^6 = 4r^4*(cos@)^2*(sin@)^2 r^2 = 4*(cos@)^2*(sin@)^2 r = 2cos@sin@ = sin(2@)

OpenStudy (anonymous):

Alternative curve sketching method, z=given (move 4x... quantity to other side) In the xz, y=0, z=x^6. In yz plane x=0, z=y^6. The shape is like a funky parabola with open bottom.